OpenStudy (anonymous):

What equation in slope intercept form represents the line that passes through the two points? (6.9, 5.9), (10.9, -2.1) Please and thank you everyone

OpenStudy (shamil98):

use the slope formula \[\huge \frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (anonymous):

Can you show me step by step on how to do that?

OpenStudy (shamil98):

to get your slope, from there input one of the coordinate pairs of (x,y) and your slope (m) into y = mx +b and solve for b after that you have all yours values for slope-intercept

OpenStudy (shamil98):

\[\huge \frac{ -2.1-5.9 }{ 10.9-6.9 } = \frac{ -8 }{ 4 } = -2\]

OpenStudy (shamil98):

\[\huge y = mx + b\] let's use (6.9,5.9) \[\huge 5.9 = 2(6.9) + b\] \[\huge 5.9 = 13.8 + b\] \[\huge -7.9 = b\]

OpenStudy (shamil98):

\[\huge y = 2x - 7.9\]

OpenStudy (anonymous):

Thank you for the steps. This has been difficult to understand for me. Thank you again.